Circular No. 8782

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006nq AND 2006ox–2006qr Six apparent supernovae have been found on unfiltered CCD images:

2006ox by L. A. G. Monard (cf. IAUC 8779); 2006qn, 2006qo, 2006qq, and 2006qr by N. Joubert, R. R. Prasad, and W. Li (LOSS/KAIT; cf. IAUC 8779); and 2006qp by K. Itagaki (cf. IAUC 8771; via S. Nakano). Also, J. Frieman and the SDSS II collaboration report (on CBET 762) the discovery of 41 new supernovae (designated 2006oy–2006qm), all in the magnitude range g = 21–23 (most are type-Ia supernovae, with 2006qk being a probable type-Ic hypernova and 2006oz a probable type-Ib supernova, along with a couple of type-II supernovae).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ox Nov. 26.75 19^{h}18^{m}24.^{s}21 −60^{o}30′09′′.9 15.3 13′′ E, 7′′ S
2006qn Nov. 29.51 7 47 44.54 +48 13 18.4 17.9 2′′.5 W, 4′′.0 S
2006qo Nov. 29.57 8 00 08.37 +56 22 07.7 18.9 1′′.7 W, 10′′.9 N
2006qp Nov. 25.85 14 42 30.65 +28 43 25.9 17.1 35′′ W, 10′′ S
2006qq Nov. 30.37 5 19 50.43 −20 58 06.4 17.0 7′′.0 E, 4′′.2 S
2006qr Nov. 30.50 8 38 01.15 − 9 49 06.1 17.9 2′′.1 W, 5′′.5 N

Additional magnitudes for 2006ox in NGC 6769: Oct. 29.75 UT, [18.0; Nov. 27.739, 15.4 ± 0.2. Monard adds that 2006ox is not present on the Digitized Sky Survey (limiting red mag 20.5). Additional KAIT magnitudes for 2006qn in UGC 4022: Nov. 16.47, [19.2; 30.46, 17.8. Additional magnitudes for 2006qo in UGC 4133 (which itself is interacting with UGC 4134): 2003 Feb. 23.23, [19.6; 2006 Nov. 19.52, [18.9; 30.47, 18.5. SN 2006qn appears to be a type-II supernova, while 2006qo is probably a type-Ia supernova (details on CBET 765). Additional magnitudes for 2006qp in NGC 5735: Aug. 15, [18.5 (Itagaki); Nov. 30.86, I ∼ 17.5 (H. Naito, Nishi-Harima Astronomical Observatory; and M. Shimada, T. Takaki, and H. Yamaoka, Kyushu University; NHAO 0.60-m reflector). Nothing is visible at the location of 2006qp on Itagaki’s images prior to Aug. 15 (limiting mag 19.0) or on the Digitized Sky Survey. Additional KAIT magnitudes for 2006qq in ESO 553-G36: Oct. 27.40, [19.0; Dec. 1.38, 16.9. Additional magnitudes for 2006qr in MCG –02-22-23: Nov. 16.52, [19.0; Dec. 1.49, 17.9. Both 2006qq and 2006qr are type-II supernovae (details on CBET 766).

T. Glinos reports that SN 2006nq (cf. IAUC 8776) appeared at V ∼ 17.9 on unfiltered images taken by D. Levy, W. Levy, and himself with a 0.64-m f/7.2 reflector at Vail, AZ, around Nov. 19.134 UT, the new object’s position measured as α = 2^{h}04^{m}24^{s}.03, δ = +9^{o}40′37′′.6 (equinox 2000.0).

2006 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green